OpenStudy (anonymous):

A store allows customers to fill their own bags of candy. Troy decides he only wants chocolate-covered pretzels and gumdrops. Chocolate- covered pretzels sell for $0.89 per pound, and gumdrops sell for $0.65 per pound. Troy's bag weighs 1.8 pounds and it cost $1.29. Write an equation in simplest form that represents the problem described here. a. 0.24p + 1.17 = 1.8 b. 0.24p + 1.17 = 1.29 c. 0.24p + 1.602 = 1.29 d. 0.24p + 1.602 = 1.8

OpenStudy (anonymous):

Answer is B.

OpenStudy (anonymous):

thank you soo much!!!!

OpenStudy (anonymous):

X+Y=1.8 x=1.8-Y 0.89X+0.65Y=1.29